OpenStudy (anonymous):

Solve and show all work. -3x+5=1/2(14x+6)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

http://clip2net.com/s/XFyV http://clip2net.com/s/XFz5 follow the links.

OpenStudy (anonymous):

My computer will not go to links

OpenStudy (anonymous):

detailed answer. and how i got them. it's just to much typing to show all of the work. if it will not go to links. just copy and open a new window and paste the links

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

\[-3x+5=\frac{1}{2}(14x+6)\] \[-3x+5=7x+3\] \[5-3=7x+3x\] \[2=10x\] \[x=\frac{2}{10}=\frac{1}{5}\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome